# Towards a Unified Copernicus Foundation Model for Earth Vision

### v2

@@ Abstract
Advances in Earth observation (EO) foundation models have unlocked the potential of big satellite data to learn generic representations from space, benefiting a wide range of downstream applications crucial to our planet. However, most existing efforts remain limited to fixed spectral sensors, focus solely on the Earth’s surface, and overlook valuable metadata beyond imagery. In this work, we take a step towards next-generation EO foundation models with three key components: 1) Copernicus-Pretrain, a massive-scale pretraining dataset that integrates 18.7M aligned images from all major Copernicus Sentinel missions, spanning from the Earth’s surface to its atmosphere; 2) Copernicus-FM, a unified foundation model capable of processing any spectral or non-spectral sensor modality using extended dynamic hypernetworks and flexible metadata encoding; and 3) Copernicus-Bench, a systematic evaluation benchmark with 15 hierarchical downstream tasks ranging from preprocessing to specialized applications for each Sentinel mission. Our dataset, model, and benchmark greatly improve the scalability, versatility, and multimodal adaptability of EO foundation models, while also creating new opportunities to connect EO, weather, and climate research. Codes, datasets and models are available at https://github.com/zhu-xlab/Copernicus-FM.
@@ 1 Introduction
Our work opens new frontiers in three directions: (1) scaling EO pretraining to unified multimodal datasets, breaking the traditional silos between surface and atmospheric observations; (2) verifying that dynamic architectures can overcome sensor heterogeneity, a longstanding challenge in multi-source remote sensing; and (3) providing the first evidence that joint cross-modal pretraining enhances performance on both surface and atmospheric tasks. Furthermore, our efforts create novel opportunities to integrate EO foundation models with weather and climate prediction systems—for instance, using compressed EO embeddings (with rich semantic information) in addition to geographical coordinates to support climate modeling. Sec. 1 gives an introduction, Sec. 2 reviews prior work, Sec. 3 introduces the Copernicus-Pretrain dataset, Sec. 4 proposes the Copernicus-FM model, Sec. 5 presents the Copernicus-Bench benchmark, Sec. 6 discusses the potential of bridging EO and climate, and Sec. 7 concludes the paper.
@@ 3.1 Data collection
First, we divide the globe into $\sim$1M $0.25\degree\times 0.25\degree$ grid cells following the coordinate mapping of ERA5. Each cell thus covers a surface area of about 28 km $\times$ 28 km, forming the basic sample unit of the dataset. With a main interest and data availability in land and its atmosphere, we apply the land mask to filter land grids with a 0.5°$\degree$ buffer around the coastline, resulting in about 393K grids.
We then use Google Earth Engine [25] to download Sentinel images for each grid cell around the anchor year 2021. For Sentinel-3/5P and DEM, we cover the whole cell with patch sizes of about $96\times 96$, $28\times 28$, and $960\times 960$ pixels for each corresponding modality. For Sentinel-3, we filter out cloudy tiles with bright pixel percentages above 20% and randomly download eight images from the year. For Sentinel-5P, we select four atmospheric variables: NO2, CO, SO2 and O3, filter out fully NaN tiles, and download 12 monthly mean images across the year for each variable. It is worth noting that for Sentinel-3 and 5P, it is possible to have fewer images available due to cloud and NaN filtering, thus the final sequence lengths can vary for different grids. For DEM, time series is not applicable, and we simply download one image for the grid.
For Sentinel-1 and 2, the data volume is too large to cover the whole grid cell and the data can be extremely redundant. Therefore, we collect images by sampling local patches within the cells. To optimize the diversity of the data, we do not directly conduct uniform sampling inside a grid, but follow a Gaussian sampling strategy [42, 62, 54] to sample locations around top-10K populated cities with a standard deviation of 50 km. We utilize this strategy to sample $\sim$1M locations, each downloading 4 seasonal $264\times 264$ image patches. We uniformly sample and download images for an additional 40K polar locations. Next, we group these local patches into the predefined $0.25\degree\times 0.25\degree$ grid cells according to their center locations. For those remaining cells that do not contain any local patches, we newly sample and download 1–2 local patches to fill in the gaps.
@@ 3.2 Dataset characteristics
In total, Copernicus-Pretrain consists of $\sim$18M images grouped into $\sim$310K $0.25\degree\times 0.25\degree$ cells, densely covering the whole land surface and near-land ocean with eight distinct Sentinel modalities. On average, each cell contains 4 local S1/S2 time series pairs (16 images each), 8 S3 images, 7/6/5/8 S5P CO/NO2/SO2/O3 images, and 1 DEM image.
Fig. 3 shows the global distribution of the full dataset, while the detailed statistics are shown in Tab. 1. The distribution and statistics of the all-modality-aligned subset and other detailed analyses can be found in the appendix.
@@ 4.1 Model architecture
Dynamic patch embedding with spectral and variable hypernetworks
@@ 4.3 Implementation details
We pretrain Copernicus-FM with ViT-Base on the Copernicus-Pretrain dataset (220K grids with all modalities available) for 100 epochs. Data augmentations include simple resized cropping and horizontal flipping. Input image sizes are as shown in Fig. 4, with patch sizes $16\times 16$ for S1/2, $8\times 8$ for S3, $4\times 4$ for S5P, and $64\times 64$ for DEM. We use Llama 3.2 [20] to encode variable names. The Fourier encoding value ranges are [1e2, 1e9] for wavelengths, [1, 1e9] for bandwidths, [1e-4, 720] for longitudes/latitudes, [1e-3, 5.1e8] for patch area, and [1, 365.25] for time. The drop probability for metadata is 0.7. The masking ratio for MIM is 0.7. We distill RGB from DINOv2 [48], and S1/2 from SoftCon [64] with loss weights of 0.1 and 0.2, respectively. More details can be found in the appendix.
@@ 4.4 Ablation studies
The main ablation results are shown in Tab. 2, where consistent improvement can be observed when gradually adding the spectral hypernetwork bandwidth, variable hypernetwork, metadata encoding, and continual distillation. Among them, the benefits of metadata encoding appear to be the most significant, especially in non-optical modalities. This highlights the importance of metadata beyond pure imagery for remote sensing applications. Detailed ablation on each metadata (location matters most), ablation of the metadata dropping ratio (higher is better), as well as other minor ablations can be found in the appendix.
@@ 5 Copernicus-Bench
Level Name Task # Images Image Size # Classes Source License L1 Cloud-S2 seg 1699/567/551 512$\times$512$\times$13 4 CloudSEN12+ [4] CC0-1.0 Cloud-S3 seg 1197/399/399 256$\times$256$\times$21 5 new CC-BY-4.0 L2 EuroSAT-S1 cls 16200/5400/5400 64$\times$64$\times$2 10 EuroSAT-SAR [65] CC-BY-4.0 EuroSAT-S2 cls 16200/5400/5400 64$\times$64$\times$13 10 EuroSAT [30] MIT BigEarthNet-S1 cls 11894/6117/5991 120$\times$120$\times$2 19 BigEarthNet v2.0 [15] CDLA-Permissive-1.0 BigEarthNet-S2 cls 11894/6117/5991 120$\times$120$\times$12 19 BigEarthNet v2.0 [15] CDLA-Permissive-1.0 LC100Cls-S3 cls 5181/1727/1727* 96$\times$96$\times$21 23 new CC-BY-4.0 DFC2020-S1 seg 3156/986/986 256$\times$256$\times$2 10 DFC2020 [28] CC-BY-4.0 DFC2020-S2 seg 3156/986/986 256$\times$256$\times$13 10 DFC2020 [28] CC-BY-4.0 LC100Seg-S3 seg 5181/1727/1727* 96$\times$96$\times$21 (288$\times$288) 23 new CC-BY-4.0 L3 Flood-S1 cd 3000/1000/1000* 224$\times$224$\times$2 3 Kuro Siwo [9] MIT LCZ-S2 cls 15000/5000/5000 32$\times$32$\times$10 17 So2Sat LCZ42 [75] CC-BY-4.0 Biomass-S3 reg 3000/1000/1000* 96$\times$96$\times$21 (288$\times$288) 1 new CC-BY-4.0 AQ-NO2-S5P reg 1480/493/494* 56$\times$56$\times$1 1 new CC-BY-4.0 AQ-O3-S5P reg 1480/493/494* 56$\times$56$\times$1 1 new CC-BY-4.0
Table 4: Benchmark results with representative single-, dual-, and multi-modal foundation models on Copernicus-Bench. We report three-run averages with standard deviations. *: patch size 16 for S1/2 (8 for S3, 4 for S5P). Best scores in bold.
@@ 5.1 Datasets in Copernicus-Bench
Copernicus-Bench consists of 15 datasets with varied Sentinel modalities organized into three application levels in practice: preprocessing, base applications, and specialized applications. Level-1 includes two cloud detection tasks from S2/3, level-2 includes 8 land use land cover classification/segmentation tasks from S1/2/3, and level-3 includes 5 specialized tasks from S1/2/3/5P. Among all the datasets, 9 are derived from existing datasets with permissive licenses, and 6 are newly curated to fill in the gap in ML-ready datasets for S3/S5P. The detailed characteristics are shown in Tab. 3. The curation process for each dataset and detailed benchmark analyses can be found in the appendix.
@@
Input Source Temperature (°C) Precipitation (m) Surface press. (Pa) Sea-level press. (Pa) $u$ wind (m/s) $v$ wind (m/s) Avg Std Avg Std Avg Std Avg Std Avg Std Avg Std Coord. (raw) 8.36 3.06 0.79 0.50 7305.91 141.38 290.68 164.84 1.12 0.55 0.95 0.58 ± $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.37$ $\pm\,0.01$ $\pm\,0.02$ $\pm\,0.01$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ Coord. (FE) 3.99 2.09 0.61 0.42 6537.54 101.09 200.93 108.12 0.96 0.51 0.80 0.54 ± $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,1.78$ $\pm\,0.18$ $\pm\,0.09$ $\pm\,0.01$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ Embed. 2.66 1.72 0.47 0.33 1627.21 97.53 197.17 117.57 0.94 0.43 0.78 0.43 ± $\pm\,0.01$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,30.13$ $\pm\,0.71$ $\pm\,0.62$ $\pm\,0.86$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ Embed. + coord. 1.98 1.35 0.46 0.32 1676.39 78.53 174.80 90.37 0.91 0.42 0.76 0.43 ± $\pm\,0.01$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,33.64$ $\pm\,0.08$ $\pm\,0.21$ $\pm\,0.15$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ Coord. (SatCLIP [36]) 3.67 1.58 0.48 0.32 3178.91 97.74 202.76 110.71 0.92 0.43 0.79 0.45 ± $\pm\,0.03$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,8.75$ $\pm\,0.19$ $\pm\,0.27$ $\pm\,0.41$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$ $\pm\,0.00$
@@ 6 Bridging EO & climate via grid embeddings
To evaluate this potential, we curate a simple set of climate tasks predicting the 10-year mean and standard deviation of 6 important climate parameters (2 m mean air temperature, annual total precipitation, surface pressure, mean sea-level pressure, and $u$/$v$ component of 10-meter wind speed) using geocoordinates or geographical representations. We randomly sample 10K grids around the globe, split them into train/val/test subsets, and calculate the corresponding climate parameters from ERA5 as targets. As a baseline, the central coordinates of the grids are sent through a linear regression model for prediction. As shown in Tab. 5, the two baseline models with raw or Fourier-encoded coordinates as input verify the basic correlation between geography and climate. We then use Copernicus-FM-derived grid embeddings as input, obtained by simply averaging model-encoded images from various modalities. These embeddings provide a richer, high-level representation of each grid’s environmental characteristics and consistently outperform coordinate-based baselines. Moreover, combining grid embeddings with coordinates yields the best overall performance, demonstrating the complementary nature of spatial context and EO-derived semantic features. Notably, even when compared against location encodings from a specialized location encoder trained on EO images [36], the grid embeddings still outperform, emphasizing the unique value of EO representations for climate studies. Looking ahead, we envision incorporating these grid embeddings into ML-based training for medium-range weather forecasting, expanding the set of static variables with EO-derived visual representations, or the set of dynamic variables with representation time series.
@@ 7 Conclusion
This work presents a series of efforts towards next-generation EO foundation models, advancing existing approaches in data, model flexibility, and benchmarking. We introduce Copernicus-Pretrain, extending existing pretraining datasets to all major Copernicus Sentinel missions, encompassing both Earth’s surface and its atmosphere. We propose Copernicus-FM, leveraging dynamic hypernetworks to process any spectral or non-spectral modality, enhancing adaptability across diverse EO sensors. We establish Copernicus-Bench, a comprehensive benchmark with hierarchical downstream tasks across various Sentinel modalities. Furthermore, we demonstrate the benefits of EO grid embeddings in a simple climate prediction task, highlighting their potential in bridging EO with weather and climate studies. This paves the way towards unified EO and climate foundation models, unlocking new possibilities for integrated Earth system understanding.
@@ A.2 Extended statistics
Fig. 9 presents the histograms of the time series lengths for S1 and S2 in the full dataset, while Fig. 10 shows the corresponding histograms in the joint subset. Similarly, Fig. 11 (left and right) presents S3 in the full dataset and joint subset, and Figs. 12 and 13 present S5P in the full dataset and joint subset.

### v3

@@ Abstract
Advances in Earth observation (EO) foundation models have unlocked the potential of big satellite data to learn generic representations from space, benefiting a wide range of downstream applications crucial to our planet. However, most existing efforts remain limited to fixed spectral sensors, focus solely on the Earth’s surface, and overlook valuable metadata beyond imagery. In this work, we take a step towards next-generation EO foundation models with three key components: 1) Copernicus-Pretrain, a massive-scale pretraining dataset that integrates 18.7M aligned images from all major Copernicus Sentinel missions, spanning from the Earth’s surface to its atmosphere; 2) Copernicus-FM, a unified foundation model capable of processing any spectral or non-spectral sensor modality using extended dynamic hypernetworks and flexible metadata encoding; and 3) Copernicus-Bench, a systematic evaluation benchmark with 15 hierarchical downstream tasks ranging from preprocessing to specialized applications for each Sentinel mission. Our dataset, model, and benchmark greatly improve the scalability, versatility, and multimodal adaptability of EO foundation models, while also creating new opportunities to connect EO, weather, and climate research. Codes at https://github.com/zhu-xlab/Copernicus-FM.
@@ 1 Introduction
Our work opens new frontiers in three directions: (1) scaling EO pretraining to unified multimodal datasets, breaking the traditional silos between surface and atmospheric observations; (2) verifying that dynamic architectures can overcome sensor heterogeneity, a longstanding challenge in multi-source remote sensing; and (3) providing the first evidence that joint cross-modal pretraining enhances performance on both surface and atmospheric tasks. Furthermore, our efforts create novel opportunities to integrate EO foundation models with weather and climate prediction systems—for instance, using compressed EO embeddings (with rich semantic information) in addition to geographical coordinates to support climate modeling.
@@ 3.1 Data collection
First, we divide the globe into $\sim$1M $\degree\times 0.25\degree$ 0.25 ° grid cells following the coordinate mapping of ERA5. Each cell thus covers a surface area of about 28 km $\times$ 28 km, forming the basic sample unit of the dataset. With a main interest and data availability in land and its atmosphere, we apply the land mask to filter land grids with a 0.5$\degree$ buffer around the coastline, resulting in about 393K grids.
We then use Google Earth Engine [25] to download Sentinel images for each grid cell around the anchor year 2021. For Sentinel-3/5P and DEM, we cover the whole cell with patch sizes of about $\times 96$ 96, $\times 28$ 28, and $\times 960$ 960 pixels for each corresponding modality. For Sentinel-3, we filter out cloudy tiles with bright pixel percentages above 20% and randomly download eight images from the year. For Sentinel-5P, we select four atmospheric variables: NO2, CO, SO2 and O3, filter out fully NaN tiles, and download 12 monthly mean images across the year for each variable. It is worth noting that for Sentinel-3 and 5P, it is possible to have fewer images available due to cloud and NaN filtering, thus the final sequence lengths can vary for different grids. For DEM, time series is not applicable, and we simply download one image for the grid.
For Sentinel-1 and 2, the data volume is too large to cover the whole grid cell and the data can be extremely redundant. Therefore, we collect images by sampling local patches within the cells. To optimize the diversity of the data, we do not directly conduct uniform sampling inside a grid, but follow a Gaussian sampling strategy [42, 62, 54] to sample locations around top-10K populated cities with a standard deviation of 50 km. We utilize this strategy to sample $\sim$1M locations, each downloading 4 seasonal $\times 264$ 264 image patches. We uniformly sample and download images for an additional 40K polar locations. Next, we group these local patches into the predefined $\degree\times 0.25\degree$ 0.25 ° grid cells according to their center locations. For those remaining cells that do not contain any local patches, we newly sample and download 1–2 local patches to fill in the gaps.
@@ 3.2 Dataset characteristics
In total, Copernicus-Pretrain consists of $\sim$18M images grouped into $\sim$310K $\degree\times 0.25\degree$ 0.25 ° cells, densely covering the whole land surface and near-land ocean with eight distinct Sentinel modalities. On average, each cell contains 4 local S1/S2 time series pairs (16 images each), 8 S3 images, 7/6/5/8 S5P CO/NO2/SO2/O3 images, and 1 DEM image.
Fig. 3 shows the global distribution of the full dataset, while the detailed statistics are shown in Tab. 1. The distribution and statistics of the all-modality-aligned subset and other detailed analyses can be found in the appendix A.
@@ 4.1 Model architecture
Dynamic patch embedding via sensor-aware hypernetworks
@@ 4.3 Implementation details
We pretrain Copernicus-FM with ViT-Base on the Copernicus-Pretrain dataset (220K grids with all modalities available) for 100 epochs. Data augmentations include simple resized cropping and horizontal flipping. Input image sizes are as shown in Fig. 4, with patch sizes $\times 16$ 16 for S1/2, $\times 8$ 8 for S3, $\times 4$ 4 for S5P, and $\times 64$ 64 for DEM. We use Llama 3.2 [20] to encode variable names. The Fourier encoding value ranges are [1e2, 1e9] for wavelengths, [1, 1e9] for bandwidths, [1e-4, 720] for longitudes/latitudes, [1e-3, 5.1e8] for patch area, and [1, 365.25] for time. The drop probability for metadata is 0.7. The masking ratio for MIM is 0.7. We distill RGB from DINOv2 [48], and S1/2 from SoftCon [64] with loss weights of 0.1 and 0.2, respectively. More details can be found in the appendix B.
@@ 4.4 Ablation studies
The main ablation results are shown in Tab. 2, where consistent improvement can be observed when gradually adding the spectral hypernetwork bandwidth, variable hypernetwork, metadata encoding, and continual distillation. Among them, the benefits of metadata encoding appear to be the most significant, especially in non-optical modalities. This highlights the importance of metadata beyond pure imagery for remote sensing applications. Subtractive ablation, ablation on each metadata (location matters), the metadata dropping ratio (higher is better), as well as other minor ablations can be found in the appendix B.
@@ 5 Copernicus-Bench
Level Name Task # Images Image Size # Classes Source License L1 Cloud-S2 seg 1699/567/551 512$\times$512$\times$13 4 CloudSEN12+ [4] CC0-1.0 Cloud-S3 seg 1197/399/399 256$\times$256$\times$21 5 new CC-BY-4.0 L2 EuroSAT-S1 cls 16200/5400/5400 64$\times$64$\times$2 10 EuroSAT-SAR [65] CC-BY-4.0 EuroSAT-S2 cls 16200/5400/5400 64$\times$64$\times$13 10 EuroSAT [30] MIT BigEarthNet-S1 cls 11894/6117/5991 120$\times$120$\times$2 19 BigEarthNet v2.0 [15] CDLA-Permissive-1.0 BigEarthNet-S2 cls 11894/6117/5991 120$\times$120$\times$12 19 BigEarthNet v2.0 [15] CDLA-Permissive-1.0 LC100Cls-S3 cls 5181/1727/1727* 96$\times$96$\times$21 23 new CC-BY-4.0 DFC2020-S1 seg 3156/986/986 256$\times$256$\times$2 10 DFC2020 [28] CC-BY-4.0 DFC2020-S2 seg 3156/986/986 256$\times$256$\times$13 10 DFC2020 [28] CC-BY-4.0 LC100Seg-S3 seg 5181/1727/1727* 96$\times$96$\times$21 (288$\times$288) 23 new CC-BY-4.0 L3 Flood-S1 cd 3000/1000/1000* 224$\times$224$\times$2 3 Kuro Siwo [9] MIT LCZ-S2 cls 15000/5000/5000 32$\times$32$\times$10 17 So2Sat LCZ42 [75] CC-BY-4.0 Biomass-S3 reg 3000/1000/1000* 96$\times$96$\times$21 (288$\times$288) - new CC-BY-4.0 AQ-NO2-S5P reg 1480/493/494* 56$\times$56$\times$1 - new CC-BY-4.0 AQ-O3-S5P reg 1480/493/494* 56$\times$56$\times$1 - new CC-BY-4.0
Table 4: Benchmark results with representative single-, dual-, and multi-modal foundation models on Copernicus-Bench. We report three-run averages with standard deviations. *: patch size 16 for S1/2 (8 for S3, 4 for S5P). Random: encoder random init. Best scores in bold.
@@ 5.1 Datasets in Copernicus-Bench
Copernicus-Bench consists of 15 datasets with varied Sentinel modalities organized into three application levels in practice: preprocessing, base applications, and specialized applications. Level-1 includes two cloud detection tasks from S2/3, level-2 includes 8 land use land cover classification/segmentation tasks from S1/2/3, and level-3 includes 5 specialized tasks from S1/2/3/5P. Among all the datasets, 9 are derived from existing datasets with permissive licenses, and 6 are newly curated to fill in the gap in ML-ready datasets for S3/S5P. The detailed characteristics are shown in Tab. 3. The curation process for each dataset and detailed benchmark analyses can be found in the appendix C.
@@
[…]
@@ 6 Bridging EO & climate via grid embeddings
To evaluate this potential, we curate a simple set of climate tasks predicting the 10-year mean and standard deviation of 6 important climate parameters (2 m mean air temperature, annual total precipitation, surface pressure, mean sea-level pressure, and $u$/$v$ component of 10-meter wind speed) using geocoordinates or geographical representations. We randomly sample 10K grids around the globe, split them into train/val/test subsets, and calculate the corresponding climate parameters from ERA5 as targets. As a baseline, the central coordinates of the grids are sent through a linear regression model for prediction. As shown in Tab. 5, the two baseline models with raw or Fourier-encoded coordinates as input verify the basic correlation between geography and climate. We then use Copernicus-FM-derived grid embeddings as input, obtained by simply averaging model-encoded images from various modalities. These embeddings provide a richer, high-level representation of each grid’s environmental characteristics and consistently outperform coordinate-based baselines. Moreover, combining grid embeddings with coordinates yields the best overall performance, demonstrating the complementary nature of spatial context and EO-derived semantic features. Notably, even when compared against location encodings from a specialized location encoder trained on EO images [36], the grid embeddings still outperform, emphasizing the unique value of EO representations for climate studies. We extend a global embedding dataset at 0.25$\degree$ based on these promising findings (Fig. 6, see appendix D for details). Looking ahead, we envision incorporating these grid embeddings into ML-based training for medium-range weather forecasting, expanding the set of static variables with EO-derived visual representations, or dynamic variables with representation time series.
Figure 6: Visualization of the Copernicus-Embed-025deg dataset as a global embedding map (PCA to 3-dim).
@@ 7 Conclusion
This work presents a series of efforts towards next-generation EO foundation models, advancing existing approaches in data, model, and benchmark. We introduce Copernicus-Pretrain, extending EO pretraining datasets to all major Sentinel missions, encompassing both Earth’s surface and its atmosphere. We propose Copernicus-FM, leveraging dynamic hypernetworks to process any spectral or non-spectral modality, enhancing adaptability across diverse EO sensors. We establish Copernicus-Bench, a comprehensive benchmark with hierarchical downstream tasks across various Sentinel modalities. Furthermore, we demonstrate the benefits of EO grid embeddings in a simple climate prediction task, highlighting their potential in bridging EO with weather and climate studies. This paves the way towards unified EO and climate foundation models, unlocking new possibilities for integrated Earth system understanding in future work.
As a limitation, the scope of this work is restricted to the Sentinel series and a temporal range of about 1 year. Future work will thus also explore efficient scaling for satellite expansion, as well as native model support for multimodal fusion and time series processing.
@@ A.2 Extended statistics
Fig. 10 presents the histograms of the time series lengths for S1 and S2 in the full dataset, while Fig. 11 shows the corresponding histograms in the joint subset. Similarly, Fig. 12 (left and right) presents S3 in the full dataset and joint subset, and Figs. 13 and 14 present S5P in the full dataset and joint subset.